# A von Neumann algebraic approach to Quantum Theory on curved spacetime

### v3

@@ 3.1. VN-STDFM is a tensorial category
As can be seen from the following result, the von Neumann algebra tensor product is a tensor structure on the category VN-STDFM.
Proposition 3.4 (Prop 2a, [13]).
@@
As can be seen from [13, Prop 2b], the above result even holds for infinite tensor products. This tensor structure even yields a prescription for combining a priori given fns weights. These facts may be deduced from [13, Prop a, b, g].
Proposition 3.5 (Prop 3g, [13]).
@@
  • $(\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}})^{\prime}=(\nu_{{\mathcal{M}}})^{\prime}\otimes(\nu_{{\mathcal{N}}})^{\prime}$.
@@
3.2. The structure of VN-STDFM allows for entanglement
@@ 4. Killing local flows on compacta of Lorentzian manifolds
The starting point of our analysis is essentially a localisation of Proposition 2.3(b) of [9].
@@
  • The manifold is causal and the causally convex hull of any compact subset is compact. [28, Proposition 2.5]
Both imply that the causally convex hull of any relatively compact set is relatively compact. [37, Page 4]
@@
A vector field $Z$ on a semi-Riemannian manifold is a Killing vector field iff the stages $\Psi_{s}^{Z}$ of the local flow are isometries. ([40, 9.21])
@@
Let $Z$ be an arbitrary smooth vector field and $K$ a compact neighbourhood. Then there exists some $t_{K}>0$ such that $K$ is in the domain of each $\Psi_{s}$ with $0<|s|\leq t_{K}$. Given any relatively compact causally compatible globally hyperbolic open subset $\mathcal{O}$ there exists some $t_{\mathcal{O}}>0$ such that $\mathcal{O}$ is in the domain of each $\Psi_{s}$ with $0<|s|\leq t_{\mathcal{O}}$.
@@ 5. Killing flows on local algebras: an illustrative example based on the Klein-Gordon equation
$PG(f\circ\Psi^{-1}_{ext})_{|\Psi(\mathcal{O})}\circ\Psi=(f\circ\Psi^{-1}_{ext})_{|\Psi(\mathcal{O})}\circ\Psi=f.$
@@
    For the local algebra described in part (1) of [31, Theorem 3.4] the prescription $f\to f\circ\Psi^{-1}$ where $f\in C^{\infty}_{0}(\mathcal{O})$ lifts to a *-isomorphism $\pi_{\Psi}$ from ${\mathcal{A}}(\mathcal{O})$ to ${\mathcal{A}}(\Psi(\mathcal{O}))$ which sends $W(f)$ where $f\in C^{\infty}_{0}(\mathcal{O})$ to $W(f\circ\Psi^{-1})$.
@@
Claim 1: Firstly note that [24, Prop 5.19] suggests that if $d\mu$ is the measure coming from the volume form, then on $\mathcal{O}$ the stage $\Psi_{s}$ is measure preserving. So for any $f,g\in C_{0}^{\infty}(M)$ with support in $\mathcal{O}$ we have that $\int_{\mathcal{O}}f\overline{g}\,d\mu=\int_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})(\overline{g\circ\Psi^{-1}})\,d\mu$ where we have used the fact that $f\circ\Psi^{-1}$ and $g\circ\Psi^{-1}$ are elements of $C_{0}^{\infty}(M)$ supported on $\Psi(\mathcal{O})$. The claim now follows from [3, Theorem 4.2.9].
@@
If $(M,g)$ is connected then all local diffeomorphisms either preserve or reverse time-orientation. See the discussion preceding [40, Cor 7.9].
@@
    On passing to the Fock representation of these Weyl algebras, we note that for any $a\in{\mathcal{A}}(\mathcal{O})$ we have that $\pi_{s}(Z)(a)$ strongly converges to $a$ as $s\to 0$.
@@ 6. Quantum Killing Lie derivatives
We shall here pass to the von Neumann algebra setting assuming all our von Neumann local algebras ${\mathcal{M}}(\mathcal{O})$ to be the $\sigma$-weak closure of the matching $C^{*}$-algebras ${\mathcal{A}}(\mathcal{O})$ in some appropriate representation. (We shall elaborate more on what we mean by “appropriate representations” in the next part of the paper.) In the notation of Corollary 5.6 we shall also assume that for any $a\in{\mathcal{M}}(\mathcal{O})$ we have that $\pi_{s}^{Z}(a)$ strongly converges to $a$ as $s\to 0$. To see that this holds in the Fock space case, notice that by [14, Theorem 6] the Bogoliubov transformations which are used to construct the $\pi^{Z}_{t}$s are implemented by a unitary. Thus the $\pi_{t}^{Z}$s admit a canonical extension to the double commutants of the ${\mathcal{A}}(\mathcal{O})$s. In addition since each of the $\pi_{t}^{Z}$s described in Corollary 5.6 is contractive, the convergence of $(\pi_{t}^{Z}(a))$ noted there is actually in the $\sigma$-strong topology. By convexity each ${\mathcal{A}}(\mathcal{O})$ is $\sigma$-strongly dense in ${\mathcal{M}}(\mathcal{O})$, and hence in this case the convergence we are assuming is just a continuous extension of the convergence noted in part (2) of Corollary 5.6. We pass to considering the consequences of strong continuity at 0. Formally Quantum Killing Lie derivatives are the infinitesimal generators of the quantum local flows of Killing vector fields. These quantum Lie derivatives are defined as follows:
@@
Some technical background is required for us to get positive results.
@@ 6.1. Generators of one-parameter local groups of contractions
A straightforward modification of the proof of [41, Corollary 2.3] suffices to prove the result. Continuity at $t=0$ follows from the definition. For $t\in(-\alpha,\alpha)$ we may simply replace $x$ with $T_{t}(x)$ and apply the claims in the definition to the limit $\lim_{s\to 0}T_{s}(T_{t}(x))$. The endpoints need to be handled with more care as one needs to check for one-sided continuity. If for example $t=-\alpha$ then for $s>0$ small enough we will for any $\xi\in H$ have that $\|T_{-\alpha}(x)\xi-T_{-\alpha+s}(x)\xi\|\leq\|T_{-\alpha}\|.\|x(\xi)-T_{s}(x)(\xi)\|\to 0$ as $s\searrow 0$. ∎
@@
    $\lim_{h\searrow 0}\frac{1}{h}\int_{t}^{t+h}T_{s}(x)\,ds=\lim_{h\nearrow 0}\int_{t+h}^{t}T_{s}(x)\,ds=T_{t}(x)$.
@@
    If $\mathcal{M}$} is a factor and if there exists a normal state $\omega$ such that $\omega\circ\delta=0$, then $\delta$ is $\sigma$-weakly closable.
@@
For the second claim note that since $\overline{S(\delta)}^{w*}$ is a weak* closed ideal, there must exists a central projection $e$ such that $e{\mathcal{M}}=\overline{S(\delta)}^{w*}$. But since $\mathcal{M}$} is a factor we must have $e\in\{0,{\mathds{1}}\}$. Note that by definition $S(\delta)\subseteq\overline{\mathrm{ran}(\delta)}^{w*}$. (Here we used the fact that by convexity the $\sigma$-weak and $\sigma$-strong* closures of $\mathrm{ran}(\delta)$ agree.) So if there is indeed a normal state such that $\omega\circ\delta=0$, then that state annihilates $\overline{S(\delta)}^{w*}$ which in turn means that $e=0$, or equivalently $S(\delta)=\{0\}$. ∎
@@ 6.2. Existence of Quantum Killing Lie derivatives
Let $(M,g)$ be a globally hyperbolic Lorentzian manifold and let ${\mathcal{M}}(M)$ be as before. If for each Killing vector field $Z$ and each $\mathcal{O}\in\mathcal{K}(M,g)$ there exists a normal state $\omega_{\mathcal{O}}$ on ${\mathcal{M}}(\mathcal{O})$ for which we have that $\omega_{\mathcal{O}}\circ\delta_{Z}=0$, the operator $\delta_{Z}$ is then a $\sigma$-strong* densely defined closable unital *-derivation.
@@ 7. The selection of appropriate von Neumann algebraic representations
[19, Proposition 3.1] For any $\mathcal{O}\in\mathcal{K}(M,g)$, the operator $i\widetilde{G}_{\mathcal{O}}$ is a densely defined symmetric operator on $L^{2}(\mathcal{O},dV)$ with domain $\mathscr{D}(\mathcal{O},E)$ which extends to a bounded operator $i\overline{G}_{\mathcal{O}}$ on $L^{2}(\mathcal{O},dV)$.
@@
$|\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}g\rangle\,dV|^{2}$ $=$ $\left|\int_{M}\langle f+\mathrm{ker}(\overline{G}_{\mathcal{O}}),\overline{G}_{\mathcal{O}}(g)\rangle\,dV\right|$
@@
We will extract an inner product satisfying equation (5.30) of [30] from the above and use that to construct a quasi-free state. To do this we will make use of the correspondences described in [30, Proposition 5.2.12]. In our setting the space $\mathrm{SOL}(\mathcal{O})$ referred to by [30] will just be $\mathrm{SOL}(\mathcal{O})=\{\widetilde{G}_{\mathcal{O}}(f)\colon f\in\mathscr{D}(\mathcal{O},E)\}$ and the space $\mathcal{E}(\mathcal{O})$ the quotient space $\mathcal{E}=\mathscr{D}(\mathcal{O},E)/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ (to see this consider equation (5.3) of [30] alongside the definition of this space). If one compares equations (5.4) and (5.12) in [30], the the bilinear form $\tau$ they define on $\mathrm{SOL}$ is just $\tau(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))=\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV$. Writing $[f]$ for the equivalence class $f+\overline{\mathrm{ker}(\widetilde{G}_{\mathcal{O}})}$, it is now an exercise to see that $\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV=\int_{M}\langle x_{f},\widetilde{G}_{\mathcal{O}}(y_{g})\rangle\,dV$ for any $x_{f}\in[f]$, $y_{g}\in[g]$. Similarly $\widetilde{G}_{\mathcal{O}}f=\overline{G}_{\mathcal{O}}[f]$. That means that we may write $\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV=\int_{M}\langle[f],\overline{G}_{\mathcal{O}}[g]\rangle\,dV$. For any $f,g\in\mathscr{D}(\mathcal{O},E)$, it follows from the previously displayed inequality that
@@
What we need in order to have access to [30, Propositions 5.2.23 & 5.2.24] is much the same thing, but for $\mathrm{SOL}(\mathcal{O})$ not $\mathcal{E}(\mathcal{O})$. Since these two spaces are bijective by [30, Proposition 5.2.12] we may now define a matching inner product $\widetilde{\mu}^{\mathcal{O}}$ on $\mathrm{SOL}(\mathcal{O})$ by simply setting $\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))=\mu^{\mathcal{O}}([f],[g])$ for all $f],[g]\in\mathcal{E}$. That will then furnish us with an inner product which satisfies (5.30) of [30]. By definition (see also Theorem 5.2.24 in [30]) there is then a unique quasi-free state $\omega_{\mathcal{O}}$ corresponding to the two-point functions $(\omega_{\mathcal{O}})_{2}(f,g)=\frac{1}{4}\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)+i\tau(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))$. This follows, see [36], from the recipe for a quasi-free state described there, namely

### v4

@@ 3.1. VN-SF and VN-MA-SF are tensorial categories
As can be seen from the following result, the von Neumann algebra tensor product is a tensor structure on the category VN-SF.
Proposition 3.7 (Proposition 2a, [13]).
@@
The above result even holds for infinite tensor products (see [13, Proposition 2b]). This tensor structure also behaves well with regard to the subcategory VN-MA-SF.
Proposition 3.8 (Proposition 3g, [13]).
@@
  • $(\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}})^{\prime}=(\nu_{{\mathcal{M}}})^{\prime}\otimes(\nu_{{\mathcal{N}}})^{\prime}$ where $\nu^{\prime}$ stands for the “opposite weight” on the commutant [52, Definition VII.1.18].
@@
3.2. The structure of both VN-SF and VN-MA-SF allow for entanglement
@@ 4. Killing local flows on compacta of Lorentzian manifolds
It is known that the local algebraic formalism as proposed in [9] behaves well with respect to groups of isometries (see [9, Proposition 2.3(b)]). However Killing vector fields generally correspond to local flows of isometries. So to fully understand how concepts like Killing vector fields may be lifted to the algebraic context, we need to see to what extent the content of [9, Proposition 2.3(b)] may be recovered when dealing with local flows of isometries. The primary challenge we need to overcome is the fact that at the manifold level we are now dealing with local rather than global isometries which only exhibit “group-like” behaviour for short times. This necessitates the creation of new proof strategies.
@@
  • The manifold is causal and the causally convex hull of any compact subset is compact. [30, Proposition 2.5]
Both imply that the causally convex hull of any relatively compact set is relatively compact. [40, Page 4]
@@
A vector field $Z$ on a semi-Riemannian manifold is a Killing vector field iff the stages $\Psi_{s}^{Z}$ of the local flow are isometries. ([43, 9.21])
@@
Let $Z$ be an arbitrary smooth vector field and $K$ a compact neighbourhood. Then there exists some $t_{K}>0$ such that $K$ is in the domain of each $\Psi^{Z}_{s}$ with $0<|s|\leq t_{K}$. Given any relatively compact causally compatible globally hyperbolic open subset $\mathcal{O}$ there exists some $t_{\mathcal{O}}>0$ such that $\mathcal{O}$ is in the domain of each $\Psi^{Z}_{s}$ with $0<|s|\leq t_{\mathcal{O}}$.
@@ 5. Killing flows on local algebras: an illustrative example based on the Klein-Gordon equation
$PG(f\circ\Psi^{-1}_{ext}){\upharpoonright}{\Psi(\mathcal{O})}\circ\Psi=(f\circ\Psi^{-1}_{ext}){\upharpoonright}{\Psi(\mathcal{O})}\circ\Psi=f.$
@@
    For the local algebra described in part (1) of [34, Theorem 3.4] the prescription $f\to f\circ\Psi^{-1}$ where $f\in C^{\infty}_{0}(\mathcal{O})$ lifts to a *-isomorphism $\pi_{\Psi}$ from ${\mathcal{A}}(\mathcal{O})$ to ${\mathcal{A}}(\Psi(\mathcal{O}))$ which sends $W(f)$ where $f\in C^{\infty}_{0}(\mathcal{O})$ to $W(f\circ\Psi^{-1})$.
@@
Claim 1: Firstly note that [25, Prop 5.19] suggests that if $d\mu$ is the measure coming from the volume form, then on $\mathcal{O}$ the stage $\Psi_{s}$ is measure preserving. So for any $f,g\in C_{0}^{\infty}(M)$ with support in $\mathcal{O}$ we have that $\int_{\mathcal{O}}f\overline{g}\,d\mu=\int_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})(\overline{g\circ\Psi^{-1}})\,d\mu$ where we have used the fact that $f\circ\Psi^{-1}$ and $g\circ\Psi^{-1}$ are elements of $C_{0}^{\infty}(M)$ supported on $\Psi(\mathcal{O})$. The claim now follows from [3, Theorem 4.2.9].
@@
If $(M,g)$ is connected then all local diffeomorphisms either preserve or reverse time-orientation. See the discussion preceding [43, Cor 7.9].
@@
    On passing to the Fock representation of these Weyl algebras one typically encounters in specific examples (see for example [15], [3, §4.7] and [32, 5.2.24]), we note that for any $a\in{\mathcal{A}}(\mathcal{O})$ we have that $\pi_{s}(Z)(a)$ strongly converges to $a$ as $s\to 0$.
@@ 6. Quantum Killing Lie derivatives
We shall here pass to the von Neumann algebra setting assuming all our von Neumann local algebras ${\mathcal{M}}(\mathcal{O})$ to be the $\sigma$-weak closure of the matching $C^{*}$-algebras ${\mathcal{A}}(\mathcal{O})$ in some appropriate representation. (We shall elaborate more on what we mean by “appropriate representations” in the next part of the paper.) In the notation of Corollary 5.6 we shall also assume that for any $a\in{\mathcal{M}}(\mathcal{O})$ we have that $\pi_{s}^{Z}(a)$ strongly converges to $a$ as $s\to 0$. To see that this holds in the Fock space case alluded to in Corollary 5.6, notice that by [14, Theorem 6] the Bogoliubov transformations which are used to construct the $\pi^{Z}_{t}$s are implemented by a unitary. Thus the $\pi_{t}^{Z}$s admit a canonical extension to the double commutants of the ${\mathcal{A}}(\mathcal{O})$s. In addition since each of the $\pi_{t}^{Z}$s described in Corollary 5.6 is contractive, the convergence of $(\pi_{t}^{Z}(a))$ noted there is actually in the $\sigma$-strong topology. By convexity each ${\mathcal{A}}(\mathcal{O})$ is $\sigma$-strongly dense in ${\mathcal{M}}(\mathcal{O})$, and hence in this case the convergence we are assuming is just a continuous extension of the convergence noted in part (2) of Corollary 5.6. We pass to considering the consequences of strong continuity at 0. Formally Quantum Killing Lie derivatives are the infinitesimal generators of the quantum local flows of Killing vector fields. These quantum Lie derivatives are defined as follows:
@@
There is a very well-developed theory of semigroups of operators which we may use as a guide in our development. However we are here not dealing with semigroups and there is no obvious way to reduce the questions we wish to answer to the semigroup setting. The tools we shall need in our analysis therefore need to be developed from scratch. We proceed with developing the technical background we require to get positive results.
@@ 6.1. Generators of one-parameter local groups of contractions
A straightforward modification of the proof of [44, Corollary 2.3] suffices to prove the result. Continuity at $t=0$ follows from the definition. For $t\in(-\alpha,\alpha)$ we may simply replace $x$ with $T_{t}(x)$ and apply the claims in the definition to the limit $\lim_{s\to 0}T_{s}(T_{t}(x))$. The endpoints need to be handled with more care as one needs to check for one-sided continuity. If for example $t=-\alpha$ then for $s>0$ small enough we will for any $\xi\in H$ have that $\|T_{-\alpha}(x)\xi-T_{-\alpha+s}(x)\xi\|\leq\|T_{-\alpha}\|.\|x(\xi)-T_{s}(x)(\xi)\|\to 0$ as $s\searrow 0$. ∎
@@
    $\lim_{h\searrow 0}\frac{1}{h}\int_{t}^{t+h}T_{s}(x)\,ds=\lim_{h\nearrow 0}\int_{t-h}^{t}T_{s}(x)\,ds=T_{t}(x)$.
@@
    If there exists a normal state $\omega$ which is faithful on the centre of $\mathcal{M}$} and for which we have that $\omega\circ\delta=0$, then $\delta$ is $\sigma$-weakly closable.
@@
For the second claim note that since $\overline{S(\delta)}^{w*}$ is a weak* closed ideal, there must exists a central projection $e$ such that $e{\mathcal{M}}=\overline{S(\delta)}^{w*}$. Note that by definition $S(\delta)\subseteq\overline{\mathrm{ran}(\delta)}^{w*}$. (Here we used the fact that by convexity the $\sigma$-weak and $\sigma$-strong* closures of $\mathrm{ran}(\delta)$ agree.) So if there is indeed a normal state such that $\omega\circ\delta=0$, then that state annihilates $\overline{S(\delta)}^{w*}$. Faithfulness of $\delta$ on the centre would then ensure that $e=0$, which then forces $S(\delta)=\{0\}$. ∎
@@ 6.2. Existence of Quantum Killing Lie derivatives
Let $(M,g)$ be a globally hyperbolic Lorentzian manifold and let ${\mathcal{M}}(M)$ be as before. If for each Killing vector field $Z$ and each $\mathcal{O}\in\mathcal{K}(M,g)$ there exists a normal state $\omega_{\mathcal{O}}$ on ${\mathcal{M}}(\mathcal{O})$ which is faithful on the centre of ${\mathcal{M}}(\mathcal{O})$ and for which we have that $\omega_{\mathcal{O}}\circ\delta_{Z}=0$, the operator $\delta_{Z}$ is then a $\sigma$-strong* densely defined closable unital *-derivation.
@@ 7. The selection of appropriate von Neumann algebraic representations
[20, Proposition 3.1] For any $\mathcal{O}\in\mathcal{K}(M,g)$, the operator $i\widetilde{G}_{\mathcal{O}}$ is a densely defined symmetric operator on $L^{2}(\mathcal{O},dV)$ with domain $\mathscr{D}(\mathcal{O},E)$ which extends to a bounded operator $i\overline{G}_{\mathcal{O}}$ on $L^{2}(\mathcal{O},dV)$.
@@
$\left|\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}g\rangle\,dV\right|$ $=$ $\left|\int_{M}\langle f+\mathrm{ker}(\overline{G}_{\mathcal{O}}),\overline{G}_{\mathcal{O}}(g)\rangle\,dV\right|$
@@
We will extract an inner product satisfying equation (5.30) of [32] from the above and use that to construct a quasi-free state. To do this we will make use of the correspondences described in [32, Proposition 5.2.12]. In our setting the space $\mathrm{SOL}(\mathcal{O})$ referred to by [32] will just be $\mathrm{SOL}(\mathcal{O})=\{\widetilde{G}_{\mathcal{O}}(f)\colon f\in\mathscr{D}(\mathcal{O},E)\}$ and the space $\mathcal{E}(\mathcal{O})$ the quotient space $\mathcal{E}({\mathcal{O}})=\mathscr{D}(\mathcal{O},E)/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ (to see this consider equation (5.3) of [32] alongside the definition of this space). If one compares equations (5.4) and (5.12) in [32], the the bilinear form $\tau$ they define on $\mathrm{SOL}$ is just $\tau(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))=\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV$. Writing $[f]$ for the equivalence class $f+\overline{\mathrm{ker}(\widetilde{G}_{\mathcal{O}})}$, it is now an exercise to see that $\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV=\int_{M}\langle x_{f},\widetilde{G}_{\mathcal{O}}(y_{g})\rangle\,dV$ for any $x_{f}\in[f]$, $y_{g}\in[g]$. Similarly $\widetilde{G}_{\mathcal{O}}f=\overline{G}_{\mathcal{O}}[f]$. That means that we may write $\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV=\int_{M}\langle[f],\overline{G}_{\mathcal{O}}[g]\rangle\,dV$. For any $f,g\in\mathscr{D}(\mathcal{O},E)$, it follows from the previously displayed inequality that
@@
What we need in order to have access to [32, Propositions 5.2.23 & 5.2.24] is much the same thing, but for $\mathrm{SOL}(\mathcal{O})$ not $\mathcal{E}(\mathcal{O})$. Since these two spaces are bijective by [32, Proposition 5.2.12] we may now define a matching inner product $\widetilde{\mu}^{\mathcal{O}}$ on $\mathrm{SOL}(\mathcal{O})$ by simply setting $\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))=\mu^{\mathcal{O}}([f],[g])$ for all $f],[g]\in\mathcal{E}$. That will then furnish us with an inner product which satisfies (5.30) of [32], namely that
(7.3) $\frac{1}{4}|\tau(x,y)|^{2}\leq\widetilde{\mu}^{\mathcal{O}}(x,x).\widetilde{\mu}^{\mathcal{O}}(y,y)\mbox{ for all }x,y\in{\mathrm{SOL}}(\mathcal{O}).$
By definition (see also Theorem 5.2.24 in [32]) there is then a unique quasi-free state $\omega_{\mathcal{O}}$ corresponding to the two-point functions $(\omega_{\mathcal{O}})_{2}(f,g)=\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)+i\frac{1}{2}\tau(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))$. This follows, see [39], from the recipe for a quasi-free state described there, namely
@@
7.1. Constructing Araki-Woods factors
We provide further details on the construction of Araki-Woods factors in the present setting. Our starting point is [32, Proposition 5.2.23 & Theorem 5.2.24]. Our exposition makes use of details of the proof of [33, Proposition 3.1]. Let $\mathcal{O}\in\mathcal{K}(M,g)$ and consider the quasi-free state on ${\mathcal{M}}(\mathcal{O})$ with the two point function given by
$(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)\to\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)+i\frac{1}{2}\tau(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)$
where $f,g\in\mathscr{D}(\mathcal{O},E)$. The completion of the vector space ${\mathrm{SOL}}(\mathcal{O})$ under the inner product $\widetilde{\mu}^{\mathcal{O}}$ then yields a real Hilbert space $H_{\mu}$. Our task is to turn this real Hilbert space into a complex Hilbert space. The first step in this endeavour is to notice that an application of the Riesz Lemma to equation (7.3) ensures there exists a contractive operator $A_{\mathcal{O}}$ on $H_{\mu}$ satisfying $A_{\mathcal{O}}^{*}=-A_{\mathcal{O}}$ such that
$\frac{1}{2}\tau(x,y)=\widetilde{\mu}^{\mathcal{O}}(x,A_{\mathcal{O}}y)\mbox{ for all }x,y\in{\mathrm{SOL}}(\mathcal{O}).$
In the present setting the non-degeneracy of $\tau$ ensures that $A_{\mathcal{O}}$ is injective. The isometry $\mathcal{J}_{{\mathcal{O}}}$ in the polar decomposition of $A_{\mathcal{O}}$ can then be shown to satisfy $\mathcal{J}_{{\mathcal{O}}},A_{\mathcal{O}}]=0$, $\mathcal{J}_{{\mathcal{O}}}^{2}=-{\mathds{1}}$ and $\mathcal{J}_{{\mathcal{O}}}^{*}=-\mathcal{J}_{{\mathcal{O}}}$. $H_{\mu}$ may now be made into a complex Hilbert space $\mathcal{H}_{\mu}$ by for all $x,y\in{\mathrm{SOL}}(\mathcal{O})$ defining a new inner product as follows:
$\langle\!\langle x,y\rangle\!\rangle_{\mathcal{O}}$ $=\widetilde{\mu}^{\mathcal{O}}(x,y)+i\widetilde{\mu}^{\mathcal{O}}(x,\mathcal{J}_{{\mathcal{O}}}y)$
$ix$ $=-\mathcal{J}_{{\mathcal{O}}}x$
We next define an operator $K_{\mathcal{O}}:{\mathrm{SOL}}(\mathcal{O})\to\mathcal{H}_{\mu}\oplus\mathcal{H}_{\mu}$ by means of the prescription
$K_{\mathcal{O}}(x)=[\tfrac{1}{2}(|A_{\mathcal{O}}|+{\mathds{1}})]^{1/2}x\oplus C[\tfrac{1}{2}({\mathds{1}}-|A_{\mathcal{O}}|)]^{1/2}x$
[…]
The approach to construct the GNS space for $\omega_{\mathcal{O}}$ described above, is the Kay-Wald approach. The Araki-Woods approach is similar and yields a copy of the above GNS space. However it uses a different approach to construct a complex inner product. We shall refer to [26, §4.8.2] to elucidate this approach. Specifically with $A_{\mathcal{O}}$ and $\mathcal{J}_{{\mathcal{O}}}$ as before this approach uses the complex inner product on ${\mathrm{SOL}}(\mathcal{O})$ given by
$\langle x,y\rangle_{AW}=-\tau(x,\mathcal{J}_{{\mathcal{O}}}y)+i\tau(x,y).$
(The sign differences in the formula recorded here when compared with that in [26] is for the purpose of aligning that two exposition with that of [33].) The two inner products are related in the sense that for all $x,y\in{\mathrm{SOL}}(\mathcal{O})$ one has
$\langle x,y\rangle_{AW}$ $=$ $-\tau(x,\mathcal{J}_{{\mathcal{O}}}y)+i\tau(x,y)$
$=$ $-2\widetilde{\mu}^{{\mathcal{O}}}(x,A_{\mathcal{O}}\mathcal{J}_{{\mathcal{O}}}y)-i\tau(x,\mathcal{J}_{{\mathcal{O}}}^{2}y)$
$=$ $2\widetilde{\mu}^{{\mathcal{O}}}(x,|A_{\mathcal{O}}|y)-i2\widetilde{\mu}^{{\mathcal{O}}}(x,\mathcal{J}_{{\mathcal{O}}}^{2}A_{\mathcal{O}}y)$
$=$ $2[\widetilde{\mu}^{{\mathcal{O}}}(x,|A_{\mathcal{O}}|y)+i\widetilde{\mu}^{{\mathcal{O}}}(x,\mathcal{J}_{{\mathcal{O}}}|A_{\mathcal{O}}|y)]$
$=$ $2\langle\!\langle x,|A_{\mathcal{O}}|y\rangle\!\rangle_{\mathcal{O}}$
[…]
In view of the above correspondence we shall where convenient pass to the Araki-Woods representation to prove some of the facts required subsequently.